An Over Damped harmonic oscillator satisfied the equation (d²x/dt²)+(10dx/dt)+16x=0. At time t=0 the particle is projected from the point x=1 towards the origin with the speed "U". Find x(t) in subsequent motion.

1
Expert's answer
2022-01-10T14:09:21-0500

The equation of motion of overdamped oscillator is given by

"\frac{d²x}{dt²}+10\frac{dx}{dt}+16x=0"

Let

"x=e^{\lambda t}"

The characteristic equation:

"\lambda^2+10\lambda+16=0"

Roots:

"\lambda_1=-8,\quad \lambda_2=-2"

Solution:

"x(t)=C_1e^{-8t}+C_2e^{-2t}"

The initial conditions

"x(0)=1,\quad x'(0)=-u"

give

"x(0)=C_1+C_2=1\\
x'(0)=-8C_1-2C_2=-u"

Finally, we get

"x(t)=\frac{1}{3}\left(u/2-1\right)e^{-8t}-\frac{1}{3}\left(u/2-4\right)e^{-2t}"